# Engineering and Computer Science, DGIST, South Korea Bridging Geometric and Semantic Foundation Models

### v1

@@ Abstract
1 Introduction
Figure 1: Qualitative comparison with DepthAnything-V1 [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao]. Our method excels in capturing more complex visual elements and details.
@@ 1 Introduction
Historically, MDE relies on handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng]. However, the advent of CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow] and ViT-based models [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Zhao et al.(2022)Zhao, Zhang, Poggi, Tosi, Guo, Zhu, Huang, Tang, and Mattoccia, Cheng et al.(2021)Cheng, Zhang, and Tang, Bae et al.(2023)Bae, Moon, and Im] has significantly enhanced its performance. Recent studies [Kendall et al.(2017)Kendall, Martirosyan, Dasgupta, Henry, Kennedy, Bachrach, and Bry, Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes, Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg, Li et al.(2023)Li, Xue, Su, He, Mao, Zhu, Sun, and Zhang, Choi et al.(2020)Choi, Jung, Lee, and Kim, Sanchez-Escobedo et al.(2018)Sanchez-Escobedo, Lin, Casas, and Pardas, Zama Ramirez et al.(2019)Zama Ramirez, Poggi, Tosi, Mattoccia, and Di Stefano, Jiao et al.(2018)Jiao, Cao, Song, and Lau, Chen et al.(2019)Chen, Liu, Liu, and Wang] have highlighted the advantages of incorporating semantic context from images, which provide more informative cues than solely relying on local geometry. More recently, MDE works [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have focused on training large-scale models using a vast amount of training datasets inspired by foundation models [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al., Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]. These approaches aim to harness the broader contextual understanding inherent in large pre-trained models, significantly advancing the field of depth estimation.
@@
Despite these advancements, current foundation models in MDE [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have primarily utilized geometric data for depth estimation and have not directly incorporated semantic information. These models require substantial training data and considerable computational resources, making it challenging to integrate additional semantic information. In response to these challenges, we propose a new integration module for MDE that adaptively incorporates semantic information, named Bridging Geometric and Semantic (BriGeS). This approach is particularly effective in addressing complex visual elements such as intricate structures, homogeneous regions, and delicate structures as shown in Fig. 1.
@@
To integrate depth features with semantic features under a resource-constrained training environment, we design Bridging Gate, which directly fuses features from the depth [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and segmentation [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] foundation model. This gate employs cross-attention and self-attention blocks to facilitate effective feature integration. To ensure high-depth estimation performance while preserving prior knowledge of a model trained on large-scale data, we exclusively train the Bridging Gate, keeping all encoders and decoders frozen. Additionally, we introduce the Attention Temperature Scaling technique to adjust the variance of the Bridging Gate’s attention map. This adjustment is crucial as the map tends to become overly concentrated on specific regions due to the fusion of two distinct modalities. By altering the distribution of attention operations during the inference stage, the technique mitigates this issue, thereby simplifying the process and enhancing overall effectiveness.
@@
2 Related Works
Monocular Depth Estimation. Monocular Depth Estimation (MDE) method estimates depth from a single image by learning the spatial shape and structure of objects. Some pioneering research in this field predominantly utilizes handcrafted features and traditional vision techniques [Hoiem et al.(2007)Hoiem, Efros, and Hebert, Liu et al.(2008)Liu, Yuen, Torralba, Sivic, and Freeman, Saxena et al.(2008)Saxena, Sun, and Ng]. However, these methods often rely solely on local information and face challenges in capturing comprehensive scene structure. To address these challenges, many MDE models have adopted CNN [Eigen et al.(2014)Eigen, Puhrsch, and Fergus, Laina et al.(2016)Laina, Rupprecht, Belagiannis, Tombari, and Navab, Ricci et al.(2018)Ricci, Ouyang, Wang, Sebe, et al., Wang et al.(2020b)Wang, Zhang, Wang, Lu, and Ruan] and ViT-based approaches [Yang et al.(2021)Yang, Tang, Ding, Sebe, and Ricci, Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun, Yuan et al.(2022)Yuan, Gu, Dai, Zhu, and Tan, Bae et al.(2023)Bae, Moon, and Im]. Recently, research has pivoted towards a self-supervised approach [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Zhou et al.(2022)Zhou, Taylor, Greenwood, and Mackiewicz, Fan et al.(2023)Fan, Tosi, Poggi, and Mattoccia, Saunders et al.(2024)Saunders, Manso, and Vogiatzis] due to the high cost of labeling training data, allowing training on larger datasets. While these models face challenges with objects having unclear boundaries or similar depths [Klingner et al.(2020)Klingner, Termöhlen, Mikolajczyk, and Fingscheidt, Dinh and Lim(2023), Shvets et al.(2024)Shvets, Zhao, Niethammer, Sengupta, and Berg], numerous studies have started integrating semantic information to refine edge predictions [Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li, Xing et al.(2023)Xing, Shen, Ho, and Tzes]. For instance, ROIFormer [Xing et al.(2023)Xing, Shen, Ho, and Tzes] proposes a semantic-aware region of interest transformer that uses geometric cues from semantic segmentation to guide the attention mechanism. TriDepth [Chen et al.(2023)Chen, Zhang, Jiang, Wang, Li, and Li] further introduces a semantic-aware triplet loss to address the edge-fattening problem, which significantly improves model performance while incurring no additional computational cost.
Zero-shot Depth Estimation. Zero-shot depth estimation involves predicting depth from images unseen during training. Earlier efforts in this domain focused on collecting more training images to enhance the generalization capabilities of MDE models [Chen et al.(2016)Chen, Fu, Yang, and Deng, Xian et al.(2018)Xian, Shen, Cao, Lu, Xiao, Li, and Luo, Chen et al.(2020)Chen, Qian, Fan, Kojima, Hamilton, and Deng]. However, data variability in unconstrained settings often leads to inconsistent depth prediction due to differing depth scales and shifts. To resolve these problems, several methods are designed to estimate affine-invariant depth, which adjusts for global scale and shift discrepancies without knowing the precise object size. A pioneering work, MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] proposes an affine-invariant loss to facilitate effective zero-shot transfer across diverse datasets by estimating relative depth. DPT [Ranftl et al.(2021)Ranftl, Bochkovskiy, and Koltun] also adapts the same loss term, and further enhances the performance of zero-shot MDE by replacing CNN-based architecture with vision transformers [Dosovitskiy(2020)]. For metric depth estimation, ZoeDepth [Bhat et al.(2023)Bhat, Birkl, Wofk, Wonka, and Müller] integrates relative and metric depth supervision, and Metric3D [Yin et al.(2023)Yin, Zhang, Chen, Cai, Yu, Wang, Chen, and Shen, Hu et al.(2024)Hu, Yin, Zhang, Cai, Long, Chen, Wang, Yu, Shen, and Shen] introduces a canonical camera space transformation module to resolve metric ambiguity. Recent works leverage the rich visual knowledge from the pre-trained model to perform zero-shot knowledge transfer. For instance, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] develops a semi-supervised training strategy using DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] encoder, while Marigold [Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] and GenPercept [Xu et al.(2024)Xu, Ge, Liu, Fan, Xie, Zhao, Chen, and Shen] use a model derived from Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer].
3 Preliminary: Foundation Models
In this section, we provide a brief overview of the foundation model used in our pipeline.
DepthAnything. The depth foundation model, DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] employs a ViT-based encoder $E_{d}$ and a CNN-based decoder $D_{d}$. The encoder processes the input image $I$ to generate features $\{f^{i}_{d}\}_{i=1}^{4}$, which serve as the input for the decoder. The decoder then utilizes this feature to regress the depth map $I_{d}$ of the input image as follows:
$\begin{gathered}I_{d}=D_{d}(\{f^{i}_{d}\}_{i=1}^{4}),~{}\{f^{i}_{d}\}_{i=1}^{4}=E_{d}(I),\\ I_{d}\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f^{i}_{d}\in\mathbb{R}^{HW\times C},~{}I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}\end{gathered}$ (1)
where $H$, $W$, and $C$ denote the height, width, and channel of the feature map, respectively, while $H^{\prime}$ and $W^{\prime}$ represent the height and width of the input image and the resulting depth.
SegmentAnything. The segmentation foundation model, SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] consists of a ViT-based encoder $E_{s}$, a prompt encoder $E_{p}$, and a decoder $D_{s}$. The encoder $E_{s}$ processes the input image $I$ to generate a feature $f_{s}$. The prompt encoder takes an input $x$, which can be points, boxes, or text, and generates a corresponding prompt token $p$. Finally, the decoder combines the feature map $f_{s}$ with the prompt token $p$ to generate the mask $M$ for the input image as follows:
$\begin{gathered}M=D_{s}(f_{s},p),~{}f_{s}=E_{s}(I),~{}p=E_{p}(x),\\ ~{}M\in\mathbb{R}^{H^{\prime}\times W^{\prime}},~{}f_{s}\in\mathbb{R}^{C\times H\times W},I\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3},\end{gathered}$ (2)
where $H^{\prime}$ and $W^{\prime}$ represent the height and width of masks.
4 Method
4.1 Overall Pipeline
@@
In this section, we describe the overall pipeline of BriGeS, illustrated in Fig. 2. Our BriGeS utilizes the encoders $E_{d}$ from DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] and $E_{s}$ from SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], along with the DepthAnything decoder $D_{d}$. We extract depth features $\{f^{i}_{d}\}_{i=1}^{4}$ and semantic feature $f_{s}$ from the respective encoders, and fuse them using Bridging Gates $\{G^{i}\}_{i=1}^{4}$ to generate semantic-aware geometric features $\mathbf{F}_{sg}=\{F^{i}_{sg}\}_{i=1}^{4}$. Since the spatial resolutions of the semantic feature $f_{s}$ and the depth feature $f_{d}^{i}$ differ, we align $f_{s}$ to match $f_{d}^{i}$ by applying bilinear interpolation followed by max pooling, resulting in the resolution-aligned semantic feature $\tilde{f}_{s}$. The same resolution-aligned semantic feature $\tilde{f}_{s}$ is shared across all Bridging Gates. The decoder then leverages these semantic-aware geometric features to regress the depth map $I^{\prime}_{d}$ of the input image as follows:
@@
Figure 2: Overview of the proposed BriGeS framework. We design a pipeline that adaptively integrates semantic information into a depth foundation model through the Bridging Gate. During inference, this integration is further refined by Attention Temperature Scaling.
4.2 Bridging Gate
While DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] demonstrates superior generalization capabilities across a wide range of unseen scenes, its lack of semantic integration can result in over-smoothed predictions in scenarios involving complex structures or ambiguous boundaries. To address this, we design the Bridging Gate that enables direct interaction between geometric and semantic information. This enhancement aims to provide more accurate depth estimations, particularly in scenes characterized by complex structures or unclear object boundaries. Our Bridging Gate ($G$) consists of a cross-attention block ($\operatorname{Block}_{c}$) and a self-attention block ($\operatorname{Block}_{s}$), designed to dynamically fuse geometric and semantic information as follows:
@@ 4.2 Bridging Gate
$\begin{gathered}F_{c}=\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right),~{}\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{d}K_{s}^{T}}{\sqrt{d}}\right)\cdot V_{s}\right),\\ Q_{d}=W_{q}^{c}(f_{d}),~{}K_{s}=W_{k}^{c}(\tilde{f}_{s}),~{}V_{s}=W_{v}^{c}(\tilde{f}_{s}),\end{gathered}$ (5)
@@
$\begin{gathered}F_{sg}=\operatorname{Block_{s}}\left(F_{c}\right),~{}\operatorname{Block_{s}}\left(F_{c}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{c}K_{c}^{T}}{\sqrt{d}}\right)\cdot V_{c}\right),\\ Q_{c}=W_{q}^{s}(F_{c}),\quad K_{c}=W_{k}^{s}(F_{c}),\quad V_{c}=W_{v}^{s}(F_{c}),\end{gathered}$ (6)
@@
4.3 Attention Temperature Scaling
@@ 4.4 Training
5 Experiments
5.1 Implementation Details
Our training methodology aligns closely with the one employed in the DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. Initially, the ground truth depth value is transformed into a disparity space and then normalized to a range of 0 to 1. To effectively train on datasets with varying depths, we employ an affine-invariant loss, which ignores the effects of scale and shift. When training BriGeS with DepthAnything-V1, only the affine-invariant loss is used, whereas training with DepthAnything-V2 incorporates both the affine-invariant loss and the gradient matching loss from MiDaS [Ranftl et al.(2020)Ranftl, Lasinger, Hafner, Schindler, and Koltun] at a 1:2 ratio.
The batch size is set to 16, with an initial learning rate of $55e-$ 5. We use the AdamW optimizer [Loshchilov(2017)] with a linear schedule. We conduct experiments on 8 RTX TITAN GPUs using labeled images, running for 2 epochs with DepthAnything-V1 and 1 epoch with DepthAnything-V2, respectively. Training our model for a single epoch takes 8 hours for DepthAnything-Base and 18 hours for DepthAnything-Large. During training, all images are resized so that the shorter side is 518 while maintaining the original aspect ratio, and then randomly cropped to a size of 518$\times$518. In addition, the temperature scaling factor $\tau$ is set to 2.5 during the inference. Our model contains 244M parameters based on the DepthAnything-Base, and 744M based on the DepthAnything-Large, including the SegmentAnything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] encoder.
5.2 Experimental Setup
Training Datasets. For training BriGeS, we employ a carefully curated subset of the extensive datasets featured in DepthAnything [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. To enhance the model’s generalization capabilities, we select three diverse datasets encompassing both indoor and outdoor environments for each model. Our training data comprises 600K samples from BlendedMVS [Yao et al.(2020)Yao, Luo, Li, Zhang, Ren, Zhou, Fang, and Quan], HRWSI [Xian et al.(2020)Xian, Zhang, Wang, Mai, Lin, and Cao], and TartanAir [Wang et al.(2020c)Wang, Zhu, Wang, Hu, Qiu, Wang, Hu, Kapoor, and Scherer] for DepthAnything-V1, and 582K samples from VKITTI [Cabon et al.(2020)Cabon, Murray, and Humenberger], Hypersim [Roberts et al.(2021)Roberts, Ramapuram, Ranjan, Kumar, Bautista, Paczan, Webb, and Susskind], and TartanAir for DepthAnything-V2, which corresponds to approximately 1% of the training datasets originally used to train the respective DepthAnything.
Method Backbone KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] ETH3D [Schops et al.(2017)Schops, Schonberger, Galliani, Sattler, Schindler, Pollefeys, and Geiger] DIODE [Vasiljevic et al.(2019)Vasiljevic, Kolkin, Zhang, Luo, Wang, Dai, Daniele, Mostajabi, Basart, Walter, et al.]

### v2

@@ Abstract
I Introduction
Figure 1: Qualitative comparison with DepthAnything-V1 [23].
@@ I Introduction
Historically, MDE relies on handcrafted features and traditional vision techniques [6]. However, the advent of CNN [4] and ViT-based models [13] has significantly enhanced its performance. Recent studies [10] have highlighted the advantages of incorporating semantic context from images, which provide more informative cues than solely relying on local geometry. More recently, MDE works [23, 24, 8] have focused on training large-scale models using a vast amount of training datasets inspired by foundation models [2, 14, 9]. These approaches aim to harness the broader contextual understanding inherent in large pre-trained models, significantly advancing the field of depth estimation.
@@
Despite these advancements, current foundation models in MDE [23, 24, 8] have primarily utilized geometric data for depth estimation and have not directly incorporated semantic information. These models require substantial training data and considerable computational resources, making it challenging to integrate additional semantic information. In response to these challenges, we propose a new integration module for MDE that adaptively incorporates semantic information, named Bridging Geometric and Semantic (BriGeS). This approach is particularly effective in addressing complex visual elements such as intricate structures, homogeneous regions, and delicate structures as shown in Fig. 1.
@@
To integrate depth features with semantic features under a resource-constrained training environment, we design Bridging Gate, which directly fuses features from the depth [23, 24] and segmentation [9] foundation model. This gate employs cross-attention and self-attention blocks to facilitate effective feature integration. To ensure high-depth estimation performance while preserving prior knowledge of a model trained on large-scale data, we exclusively train the Bridging Gate, keeping all encoders and decoders frozen. Additionally, we introduce the Attention Temperature Scaling technique to adjust the variance of the Bridging Gate’s attention map. This adjustment is crucial as the map tends to become overly concentrated on specific regions due to the fusion of two distinct modalities. By altering the distribution of attention operations during the inference stage, the technique mitigates this issue, thereby simplifying the process and enhancing overall effectiveness.
@@
II Method
II-A Overall Pipeline
@@
In this section, we describe the overall pipeline of BriGeS, illustrated in Fig. 2. Our BriGeS utilizes the encoders $E_{d}$ from DepthAnything [23, 24] and $E_{s}$ from SegmentAnything [9], along with the DepthAnything decoder $D_{d}$. We extract depth features $\{f^{i}_{d}\}_{i=1}^{4}$ and semantic feature $f_{s}$ from the respective encoders, and fuse them using Bridging Gates $\{G^{i}\}_{i=1}^{4}$ to generate semantic-aware geometric features $\mathbf{F}_{sg}=\{F^{i}_{sg}\}_{i=1}^{4}$. Since the spatial resolutions of the semantic feature $f_{s}$ and the depth feature $f_{d}^{i}$ differ, we align $f_{s}$ to match $f_{d}^{i}$ by applying bilinear interpolation followed by max pooling, resulting in the resolution-aligned semantic feature $\tilde{f}_{s}$. The same resolution-aligned semantic feature $\tilde{f}_{s}$ is shared across all Bridging Gates. The decoder then leverages these semantic-aware geometric features to regress the depth map $I^{\prime}_{d}$ of the input image as follows:
@@
II-B Bridging Gate
While DepthAnything [23, 24] demonstrates superior generalization capabilities across a wide range of unseen scenes, its lack of semantic integration can result in over-smoothed predictions in scenarios involving complex structures or ambiguous boundaries. To address this, we design the Bridging Gate that enables direct interaction between geometric and semantic information. This enhancement aims to provide more accurate depth estimations, particularly in scenes characterized by complex structures or unclear object boundaries. Our Bridging Gate ($G$) consists of a cross-attention block ($\operatorname{Block}_{c}$) and a self-attention block ($\operatorname{Block}_{s}$), designed to dynamically fuse geometric and semantic information as follows:
@@ II-B Bridging Gate
$\begin{gathered}F_{c}=\operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right),\\ \operatorname{Block}_{c}\left(f_{d},\tilde{f}_{s}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{d}K_{s}^{T}}{\sqrt{d}}\right)\cdot V_{s}\right),\\ Q_{d}=W_{q}^{c}(f_{d}),~K_{s}=W_{k}^{c}(\tilde{f}_{s}),~V_{s}=W_{v}^{c}(\tilde{f}_{s}),\end{gathered}$ (3)
@@
$\begin{gathered}F_{sg}=\operatorname{Block_{s}}\left(F_{c}\right),\\ \operatorname{Block_{s}}\left(F_{c}\right)=\operatorname{MLP}\left(\operatorname{Softmax}\left(\frac{Q_{c}K_{c}^{T}}{\sqrt{d}}\right)\cdot V_{c}\right),\\ Q_{c}=W_{q}^{s}(F_{c}),\quad K_{c}=W_{k}^{s}(F_{c}),\quad V_{c}=W_{v}^{s}(F_{c}),\end{gathered}$ (4)
@@
Figure 2: Overview of the proposed BriGeS framework.
II-C Attention Temperature Scaling
@@ II-D Training
III Experiments
TABLE I: Quantitative comparison of zero-shot relative depth estimation. (Bold: best, Underlined: second best)
Method Backbone KITTI NYUv2 ETH3D DIODE